OpenStudy (anonymous):

deidra plotted 3 points from the equation y=3x+1 on this coordinate grid.she drew a straight line through the points.which orderd pair would also be on this line (5,11) (6,18) (3,10) (5,13)

OpenStudy (anonymous):

help

jimthompson5910 (jim_thompson5910):

(5,11) means x = 5 and y = 11 x always comes first in the ordered pair. Think of x coming before y in the alphabet. What we do is plug in x = 5 and y = 11 into the equation y=3x+1. After simplifying, if we get a true equation, then (5,11) is on the line. Otherwise, (5,11) is not on the line

jimthompson5910 (jim_thompson5910):

y=3x+1 y=3*5+1 ... replace x with 5 11 = 3*5+1 ... replace y with 11 11 = 15 + 1 11 = 16 The last equation 11 = 16 is false, so (5,11) is NOT on the line y = 3x+1

jimthompson5910 (jim_thompson5910):

Do the same with the other points given

OpenStudy (anonymous):

whats the anwer

jimthompson5910 (jim_thompson5910):

try out the second point and show me your work

OpenStudy (anonymous):

(6,18)

jimthompson5910 (jim_thompson5910):

yes, plug in x = 6 and y = 18 into y=3x+1 do you get a true equation?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what equation do you get? Can I see your work?